OpenStudy (anonymous):

can someone check this for me please Write the equation of the line passing through the points (-6,1) and (-4,2). My answer: y2-y1/x2-x1 2-4/-2-0 Slope= 4 y=mx+b m=Slope B=Y intercept y=4x+4 <--- is this right?

OpenStudy (anonymous):

@Jaynator495

OpenStudy (anonymous):

@openstudier101x

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

I think you need to check the slope... (2 - 1)/(-4 - (-6)) =

OpenStudy (anonymous):

i got 0.5 @campbell_st

OpenStudy (anonymous):

so is it y=.5x+4

OpenStudy (campbell_st):

great... that's a big help... so here is a short cut... the slope intercept form is y = mx + b you know m... so its y = 0.5x + b pick either point say (-4, 2) if you substitute these into the equation you can find b, the y-intercept 2 = 0.5*(-4) + b or 2 = -2 + b what's the value of b...?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

No @Blondebeauty it is not 2

OpenStudy (anonymous):

Than i dont get it

OpenStudy (anonymous):

-2+2 is 0, not 2.

OpenStudy (campbell_st):

ok... what do you add to -2 to get and answer of 2..?

OpenStudy (anonymous):

^

OpenStudy (anonymous):

-2+2 is 0, and 0 + 2 is 2, so add those together: 2+2=4, so b is 4

OpenStudy (anonymous):

@Blondebeauty

OpenStudy (anonymous):

-2+4 equals 2

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Good job! :)

OpenStudy (anonymous):

so the equation should be 0.5x+4

OpenStudy (campbell_st):

that's correct so the y-intercept is y = 2 so b = 2 then the equation, using your information is y = 0.5x + 2

OpenStudy (anonymous):

oh okay sorry got ahead of myself

OpenStudy (campbell_st):

you can use the other point x = -6 and y = 1 so see if it's correct...

OpenStudy (anonymous):

Okay thank you

OpenStudy (anonymous):

It’s ok @Blondebeauty but yes @campbell_st is right

OpenStudy (anonymous):

...Happy studying! :)